# Overcoming Knowledge Barriers: Online Imitation Learning from Observation with Pretrained World Models

### v1

@@ 5 Related Work
Imitation Learning from Observation. ILfO [57, 58, 11, 35, 3, 65, 37] has become more popular in recent years due to their potential to utilise internet-scale videos for behaviour learning. Most of the previous works [57, 58, 35] study the problem only with the true state as observation. Recent works [11, 3, 65, 37] have started to shift toward image observations as a more general setting. Our work is a continuation of this journey.
Pretrained Models for Decision-Making. Inspired by the tremendous progress made in recent years in CV and NLP fields with the power of pretrained models, the decision-making community is also trying to follow the trend. Most recent works focus on the use of Large Language Model (LLM) for decision-making. A prompted model is used for producing trajectories and plans [10, 28, 1, 13], code [63, 36, 56, 9, 29] or for modifying the reward [39, 40]. There are also other people studying the benefit of pretrained visual models for visuomotor tasks [55, 41, 24, 44] while others try to train large policy networks directly with transformers [62] and huge datasets [8, 7, 49]. However, there is only little attention being put on pretrained world models [65, 48, 52], which are natively developed by the model-based decision-making community and perfectly fit into the pretraining and finetuning paradigm. Our work explores this overlooked domain and showcases its potential.
@@ 6 Discussion
In this paper, we identify two knowledge barriers, namely the EKB and the DKB, limiting the performance of state-of-the-art ILfO methods using pretrained models. We thoroughly analyse the underlying cause of each barrier and propose practical solutions. Specifically, we propose to use online interaction with a data-driven regulariser to solve the EKB and surrogate reward labelling to reduce the DKB. Combining these solutions, we propose AIME-v2, a method based on the recently proposed AIME and showcase its efficiency compared to state-of-the-art ILfO methods. We also conducted a series of ablation studies about how each knowledge barrier is addressed by the proposed solution and how their hyper-parameters influence the performance.
However, the proposed solutions still have drawbacks. First, the data-driven regulariser is not practical when the model is pretrained on huge datasets like the foundation models popular in the fields of CV and NLP. Reducing the amount of data needed for the regulariser could greatly improve the usability of the method. Second, although having pretrained models is beneficial, having too many pretrained components can be detrimental for model selection. Especially in AIME-v2, the world model and the VIPER model share a very similar interface. Designing a shared model that can serve both interfaces could ease the use of the method. Last but not least, due to the high demand of computing resources, we only study the pretrained world model on a very small scale. It will be an interesting direction to study these model at larger scales.
@@
Appendix A Algortihm
Algorithm 1 AIME-v2
  Input: Embodiment dataset $D_{\mathrm{body}}$, Demonstration dataset $D_{\mathrm{demo}}$, Pretrained world model parameters $\phi,\theta$, Pretrained VIPER model $p_{\nu}$, regulariser ratio $\alpha$, value gradient weight $\beta$, batch size $\mathrm{B}$
  Initialise policy and critic parameters $\psi,\xi$ randomly.
  for $i=1$ to policy pretraining iterations do
     Draw a batch of demonstrations $o_{1:T}\sim D_{\mathrm{demo}}$
     Update policy parameters $\psi$ with Equation 3.
  end for
  Initialize $D_{\mathrm{online}}\rightarrow\emptyset$.
  for $i=1$ to Environment Interaction budget do
     Collect a new episode $\{o_{1:T},a_{1:T}\}$ with the current policy $\pi_{\psi}$
     Estimate reward using VIPER $r_{1:T}^{\mathrm{VIPER}}=p_{\nu}(o_{1:T})$
     Append $\{o_{1:T},a_{1:T},r_{1:T}^{\mathrm{VIPER}}\}$ to $D_{\mathrm{online}}$
     # Update world model
     Draw $\alpha\cdot\mathrm{B}$ samples $b_{\mathrm{body}}\sim D_{\mathrm{body}}$
     Draw $(1-\alpha)\cdot\mathrm{B}$ samples $b_{\mathrm{online}}\sim D_{\mathrm{online}}$
     Define combined batch $b=b_{\mathrm{body}}\cup b_{\mathrm{online}}$
     Finetune model with batch $b$ using Equation 4.
     # Update policy
     Sample a batch from $D_{\mathrm{demo}}$
     Update policy parameters $\psi$ with Equation 8.
     Update value function parameters $\xi$ with Equation 7.
  end for
@@ Appendix B Compute Resources
All the experiments are run on a local cluster with a few A100 and RTX8000 instances. All the experiments are tuned to use less than 10GB of GPU memory so that they can run in A100 MIG. World models pretraining requires about 24 GPU hours, while VIPER models require negligible time for training. Each DMC experiment requires about 40 GPU hours while each Meta-World experiment requires about 24 GPU hours.
@@ Appendix C Hyper-parameters
Table 1: AIME-v2 hyper-parameters use for each benchmark.
DMC Meta-World
@@
CNN structure Ha and Schmidhuber [18] Hafner et al. [22]
@@
Policy entropy regulariser weight $1\mathrm{e}{-4}$
Model learning rate $3\mathrm{e}{-4}$
Policy learning rate $3\mathrm{e}{-4}$
Value network learning rate $8\mathrm{e}{-5}$
@@
Value gradient loss weight $\beta$ $0.1$
Appendix D Details for Resetting Meta-World Tasks
To generate the image observation datasets from the TD-MPC2 replay buffer [23], we modify the Meta-World codebase to reset the environment to the initial state of the trajectory from the first observation. Luckily, the starting position of the robot arm is always the same for each task, so that we do not need to apply inverse kinematics to solve for the initial pose of the robot arm. For the object and the target position, for most of the tasks, the internal reset position can be computed by making a constant shift on the object position and the target position in the observations. There are, however, also a few edge cases which we handle differently.
@@ Appendix D Details for Resetting Meta-World Tasks
For basketball and box-close, it seems like there is some internal collision detection that will alter the object and robot position after the task is reset, so computing the exact reset value from the observation is not possible. For these two tasks, we instead resort to a search-based method. To be specific, we use a gradient-free optimiser from [38] to search over the resetting space of the object and find the reset position that minimise the L2 distance with the true observation.
@@
Appendix E Overfitting of the VIPER Models
@@
Figure 5: Correlation of the VIPER reward and the real reward with models trained with different number of gradient steps. Each point represents one trajectory. We can clearly see the model gradually overfitting and losing the correlation with the real reward when training for more than 1 00010001\,0001 000 gradient steps.
To better illustrate the overfitting problem for VIPER models and justify our choice of training fewer iterations, we train the VIPER models for a varying number of gradient steps and evaluate the correlations between the VIPER reward and the true reward on both the expert dataset from PatchAIL, where the VIPER model is trained on, and the replay buffer dataset from Zhang et al. [65].
@@
Specifically, we train the same VIPER model with $\{100,500,1000,2000,5000,100000\}$ gradient steps and plot the result in Figure 5. As we can clearly see, when training with less than or equal to 1 00010001\,0001 000 gradient steps, VIPER reward has a very nice correlation with the true reward, with the middle-range performance even like a linear correlation. The best model could be selected from $500$ and $1000$ gradient steps. However, as we train the model for longer, the VIPER reward for the expert trajectories is boosted even higher, and as a side effect, it also relatively boosts up the VIPER reward for low-performance trajectories. This is because, when overfitting the expert trajectories, the model increases the marginal likelihood of all the observations in the expert trajectories to a higher value, which also includes a few frames of the robot lying on the ground at the very beginning of each trajectory after reset. For these low-performance trajectories, the robot remains mainly stuck around the initial position and struggles on the ground. This artifact of the overfitted VIPER reward creates a sharp local maximum in the low-performance region that the agent can hardly get away from.
Appendix F Additional Experiments
Figure 6: Additional ablation on DMC tasks by exploring the synergy between AIME and VIPER model. Return are calculated by running the policy 10 times with the environment and taking the average return. The results are averaged across 3 seeds with the shade region representing 95% CI.
Synergy between AIME and VIPER model. We also find there is a synergy between AIME and VIPER model. As we shown in Appendix E, one inherent problem of VIPER reward is that it not only incentivises the expert behaviour as the optimal, but also a stationary behaviour with very low reward as a local maxima. In order to work with the VIPER reward, the agent needs to have the ability to escape from the local maxima region. AIME offers the IL loss to imitate the expert demonstrations and can achieve decent performance even when pretrained offline, which helps to escape the local maxima. To better show the synergy, we provide additional ablation results with the VIPER reward in Figure 6. In the experiments, we include two other variants: the AIME-v2 w/o AIME is to remove the AIME IL loss from the online policy learning, so that the policy is pretrained by AIME loss but finetuned with only RL loss from the VIPER reward; while the VIPER is following the implementation in the original VIPER paper with RL loss on both the VIPER reward for the task and intrinsic reward for exploration. From the result, we can clearly see that without the help of AIME IL loss, VIPER reward cannot reliably motivate the agent to learn good behaviours. Even when in walker tasks, the w/o AIME variant can solve the tasks to certain extends, it depends strongly on the random seeds. In conclusion, the good performance of AIME-v2 cannot be achieved by either AIME IL loss or VIPER RL loss alone but a combination of both.

### v2

@@ 5 Related Work
Imitation Learning from Observation. ILfO (Torabi et al., 2018; 2019; DeMoss et al., 2023; Li et al., 2023; Baker et al., 2022; Zhang et al., 2023; Liu et al., 2022a) becomes more popular in recent years due to their potential to utilise internet-scale videos for behaviour learning. Most of the previous works (Torabi et al., 2018; 2019; Li et al., 2023; Kidambi et al., 2021) study the problem only with the true state as observation. Recent works (DeMoss et al., 2023; Baker et al., 2022; Zhang et al., 2023; Liu et al., 2022a) have started to shift toward image observations as a more general setting. Only few works (Zhang et al., 2023; Torabi et al., 2018) can leverage pretrained models. Our work is a continuation of this journey and further emphasise the performance benefit from pretrained models.
Pretrained Models for Decision-Making. Inspired by the tremendous progress made in recent years in CV and NLP fields with the power of pretrained models, the decision-making community is also trying to follow the trend. Most recent works focus on the use of Large Language Model (LLM) for decision-making. A prompted model is used for producing trajectories and plans (Chen et al., 2024; Huang et al., 2022; Ahn et al., 2022; Di Palo et al., 2023), code (Vemprala et al., 2023; Liang et al., 2023; Singh et al., 2022; Chen et al., 2023; Huang et al., 2023) or for modifying the reward (Ma et al., 2023; Mahmoudieh et al., 2022). There are also other people studying the benefit of pretrained visual models for visuomotor tasks (Shah & Kumar, 2021; Majumdar et al., 2023; Hansen et al., 2023b; Parisi et al., 2022) while others try to train large policy networks directly with transformers (Vaswani et al., 2017) and huge datasets (Brohan et al., 2023; ; Reed et al., 2022). However, there is only little attention being put on pretrained world models (Zhang et al., 2023; Rajeswar et al., 2023; Sekar et al., 2020), which are natively developed by the model-based decision-making community and perfectly fit into the pretraining and finetuning paradigm. Our work explores this overlooked domain and showcases its potential.
@@ 6 Discussion
In this paper, we identify two knowledge barriers, namely the EKB and the DKB, which limit the performance of state-of-the-art ILfO methods using pretrained models. We thoroughly analyse the underlying cause of each barrier and propose practical solutions. Specifically, we propose to use online interaction with a data-driven regulariser to overcome the EKB and surrogate reward labelling to reduce the DKB. Combining these solutions, we propose AIME-NoB and showcase its supreme efficiency compared to SOTA ILfO methods. Our ablation studies show how each knowledge barrier is addressed by the proposed solution and how different design choices influence the performance.
However, there are still limitations for the scope of this study. First, this work is mainly an empirical study. Some theoretical results could enhance the understanding of the knowledge barriers. Second, as suggested by the ablation study, a careful design for a schedule of $\alpha$ could further improve the sample efficiency. Third, due to the high demand of computing resources, we only study the pretrained world model on a rather small scale, i.e. the biggest model has only 20M parameters. It will be interesting to study how well the algorithm scales to larger models. Lastly, although we have shown a single pretrained world model can be used for multiple tasks, the power of pretrained world models are not fully realized. It would be interesting to see how world models can be used to train a multi-tasks policy. These limitations provide directions for future works.
@@
Acknowledgments
We would like to thank Botond Cseke for his help in reviewing the mathematical aspects of this paper.
Appendix A Implementation details of AIME-NoB variants
AIME-NoB with AIL rewards The idea of AIL (Ho & Ermon, 2016; Torabi et al., 2019) is to train a discriminator $D_{\nu}$ to classify whether the observation is from the expert demonstration or the agent’s rollout. As the observation is an image in our case, we follow previous works (Haldar et al., 2022; Liu et al., 2022a) to apply the discriminator upon the feature of an image encoder, for which we naturally reuse the image encoder $f_{\phi}$ from the world model. In this way, the discriminator is trained by
$\nu^{*}=\mathop{\arg\!\min}_{\nu}\>$ $\mathbb{E}_{o^{-}\sim D_{\mathrm{body}}\cup D_{\mathrm{online}},o^{+}\sim D_{\mathrm{demo}},\alpha\sim U(0,1)}$
$[D_{\nu}(f_{\phi}(o^{-}))-D_{\nu}(f_{\phi}(o^{+}))+\lambda||\nabla_{\nu}D_{\nu}(\alpha f_{\phi}(o^{-})+(1-\alpha)f_{\phi}(o^{+})||^{2}_{2}].$ (10)
The first two terms in Equation 10 train the discriminator to put higher score on the observations from the demonstrations while lower score on the observations generated by the agent. The last term is a regulariser to smooth out the landscape of discriminator. The regulariser weight $\lambda$ is set to 10. Following previvous works (Haldar et al., 2022; Liu et al., 2022a), the discriminator is implemented as a 3-layers MLP with 1024 hidden units for each layer and trained with Adam optimiser with learning rate of 1e-4. With a trained discriminator, the AIL reward can be computed as
$r^{\mathrm{AIL}}_{T}=\log D_{\nu}(f_{\phi}(o_{T}))-\log(1-D_{\nu}(f_{\phi}(o_{T}))).$ (11)
Since the value of AIL rewards also depends on the image encoder $f_{\phi}$, which is changing during the course of training, previous works (Liu et al., 2022a; Haldar et al., 2022) maintain a slow-update version of the image encoder to mitigate the non-stationary of the reward. But since our image encoder is pretrained as part of the world model and finetuned with the data-driven regulariser, we find its weight is stable during the course of training. Thus, we don’t apply this additional slow encoder. In the experiments, we find the adverserial training is stable with the help of the world model, as also shown in Rafailov et al. (2022).
AIME-NoB with OT rewards OT (Papagiannis & Li, 2023; Haldar et al., 2022) was introduced to imitation learning to alleviate the non-stationary reward and sensitive to hyper-parameters problems in AIL. The idea is to use the optimal transport to measure the minimal effort of moving any trajectory $\mathcal{T}=\{o_{1:T}\}$ to a demonstration trajectory $\mathcal{T}^{d}=\{o^{d}_{1:T^{d}}\}$, where the length of the trajectory $T$ and $T_{d}$ can be different. The effort of moving the trajectory is measure by a Wasserstein distance, i.e.
$g(\mu,f_{\phi}(\mathcal{T}),f_{\phi}(\mathcal{T}^{d}),c)=\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T^{d}}\mu_{t,t^{\prime}}c(f_{\phi}(o_{t}),f_{\phi}(o_{t^{\prime}}^{d})),$ (12)
where $\mu\in\mathbb{R}^{T\times T^{d}}$ is the transportation matrix that satisfy $\mu\mathbf{1}=\frac{1}{T}\mathbf{1}$ and $\mu^{T}\mathbf{1}=\frac{1}{T^{d}}\mathbf{1}$, $f_{\phi}$ is the process function that map the raw observation to a metric space for which we reuse the image encoder as in AIL, $c$ is the cost function that measure the distance between two vectors in the metric space for which we use the cosine distance. With the cost measure $g$, the optimal transport solve for the optimal transportation matrix with
$\mu^{*}=\mathop{\arg\!\min}_{\psi}g(\mu,f_{\phi}(\mathcal{T}),f_{\phi}(\mathcal{T}^{d}),c).$ (13)
With the optimal transportation matrix, the OT reward can be defined as
$r^{\mathrm{OT}}_{T}=-\lambda\sum_{t=1}^{T^{d}}\mu^{*}_{T,t}c(f_{\phi}(o_{T}),f_{\phi}(o_{t}^{d})).$ (14)
The $\lambda$ is hyper-parameter to scale the OT reward to be easier for the agent to learn with. We follow Haldar et al. (2022) to apply an adapted normalisation scheme where the scale factor is based on the cost measure of the first trajectory we evaluate, i.e.
$\lambda=\frac{4}{g(\mu^{*},f_{\phi}(\mathcal{T}^{(1)}),f_{\phi}(\mathcal{T}^{d}),c)}.$ (15)
When multiple demonstrations are available, we take the OT reward from the trajectory with the lowest total transportation cost.
It may worth note that OT is the only variant that doesn’t requires to train a separate model for the reward labelling. However, it is still changing during the course of training since the image encoder $f_{\phi}$ gets finetuned.
AIME-NoB with VIPER rewards VIPER (Escontrela et al., 2023) trains a video prediction model on the demonstration datasets and treats the likelihood of the video prediction model as the reward for policy learning, i.e.
$r^{\mathrm{VIPER}}_{T}=\log p_{\nu}(o_{T}|o_{t<T}).$ (16)
In the original paper, the authors first pretrain a VQ-GAN (Esser et al., 2021) from a multi-tasks expert dataset, and then train a GPT-style auto-regressive model in the quantised space for prediction. For a fair comparison with other variant, we consider to only train the VIPER model on the single demonstration dataset for the task. For simplicity of the implementation, in this paper, we consider training an unconditioned latent world model as in Seo et al. (2022b) to model the VIPER reward. We use the same RSSM architecture of the model learning for DMC, only removing the condition of the actions, and we train the VIPER model for each task separately. Especially during training, we find training such a powerful model from scratch on a small dataset can easily result in over-fitting. Thus, we empirically choose to train the model only for 500 gradient steps for DMC models and 1000 gradient steps for MetaWorld models. We show evidence of overfitting in Section G.1. Due to the large scale of the ELBO, we also apply symlog (Hafner et al., 2023) when computing the VIPER reward. Another difference with the original VIPER paper is that we do not use intrinsic motivation as the exploration bonus as the authors suggested, since the AIME loss for policy learning already provides task-related guidance for exploration. We only apply an entropy regulariser to the policy as is common practice. We further show the synergy between AIME and VIPER in Appendix I.
@@ Appendix B Compute Resources
All the experiments are run on a local cluster with a few A100 and RTX8000 instances. All the experiments are tuned to use less than 10GB of GPU memory so that they can run in A100 MIG. World models pretraining requires about 24 GPU hours, while VIPER models require negligible time for training. Each DMC experiment requires about 40 GPU hours while each MetaWorld experiment requires about 20 GPU hours. And for reference, the baseline algorithm AIME requires about 7 GPU hours for each task.
@@ Appendix C Hyper-parameters
Appendix D Source of Datasets
We use the expert trajectories from Haldar et al. (2022) at https://osf.io/4w69f/?view_only=e29b9dc9ea474d038d533c2245754f0c. The authors didn’t provide a License for their dataset. Further, we use the replay buffer dataset from Hansen et al. (2023a) at https://huggingface.co/datasets/nicklashansen/tdmpc2/tree/main/mt80. The authors provide the dataset under the MIT License. Moreover, we use the replay buffer dataset from Zhang et al. (2023) at https://github.com/argmax-ai/aime/tree/main/datasets. The authors provide the dataset under the CC BY 4.0 License.
Table 1: AIME-NoB hyper-parameters use for each benchmark.
DMC MetaWorld
@@ Appendix D Source of Datasets
CNN structure Ha & Schmidhuber (2018) Hafner et al. (2023)
@@
[…]
@@
Value gradient loss weight $\beta$ 1.0
Appendix E Details for Resetting MetaWorld Tasks
To generate the image observation datasets from the TD-MPC2 replay buffer (Hansen et al., 2023a), we modify the MetaWorld codebase to reset the environment to the initial state of the trajectory from the first observation. Luckily, the starting position of the robot arm is always the same for each task, so that we do not need to apply inverse kinematics to solve for the initial pose of the robot arm. For the object and the target position, for most of the tasks, the internal reset position can be computed by making a constant shift on the object position and the target position in the observations. There are, however, also a few edge cases which we handle differently.
@@ Appendix E Details for Resetting MetaWorld Tasks
For basketball and box-close, it seems like there is some internal collision detection that will alter the object and robot position after the task is reset, so computing the exact reset value from the observation is not possible. For these two tasks, we instead resort to a search-based method. To be specific, we use a gradient-free optimiser from (Liu et al., 2022b) to search over the resetting space of the object and find the reset position that minimises the L2 distance with the true observation.
@@
Appendix F Difference between compounding error and DKB
For DKB, most literature (Ho & Ermon, 2016; Peng et al., 2021; Torabi et al., 2019; Liu et al., 2022a) understand it as compounding error, in which the error is understood as when placing the policy with the same initial states as the demonstration, the learned policy will gradually diverge from the demonstrations due to error accumulation. Our DKB is a broader concept than compounding error. DKB attributes the gap of perform to the lack of learning signal on unsupported space of the demonstrations. In this way, EKB can explain more empirical successes in the literature than compounding error can. For example, a recent work MAHALO (Li et al., 2023) shows evidence of the importance of the size of the covered space. The authors studied a similar ILfO setup with embodiment and demonstration datasets. They compared two variants: for one they train an Inverse Dynamics Model (IDM) from the embodiment dataset and use it to label the demonstration dataset, while for another they train a reward model from the demonstration dataset by labelling all time steps with a reward of 1, and then use it to label the embodiment dataset. Finally, they run the same offline RL algorithm on both labelled datasets. The results show the second variant attains a much better performance even though the labelling from the reward model is not as meaningful as the actions from the IDM.
Appendix G Landscape of the surrogate rewards
To better understand the different types of surrogate reward and why one works better than the others, we investigate the relevance between the surrogate reward and the true reward. We take the one seed of the final model from each of the variant on walker-run task and evaluate the surrogate rewards on both the expert dataset from PatchAIL, where the surrogate reward model is based on, and the replay buffer dataset from Zhang et al. (2023).
Figure 7: Correlation of surrogate rewards and true rewards on the walker-run task.
As we can see from the results in Figure 7, all the surrogate rewards manage to put the expert demonstrations with a higher reward than the trajectories in the replay buffer. Among them, AIL, which performs the best, has a more linear correlation with the true reward and a higher slow for the linear regression.
G.1 Overfitting of the VIPER model
@@
Figure 8: Correlation of the VIPER reward and the real reward with models trained with different numbers of gradient steps. Each point represents one trajectory. We can clearly see the model gradually overfitting and losing the correlation with the real reward when training for more than 1000 gradient steps.
@@
To better illustrate the overfitting problem for VIPER models and justify our choice of training fewer iterations, we train the VIPER models for a varying number of gradient steps and evaluate the correlations between the VIPER reward and the true reward. Specifically, we train the same VIPER model with {100, 500, 1000, 2000, 5000, 100000} gradient steps and plot the result in Figure 8. As we can clearly see, when training with less than or equal to 1000 gradient steps, VIPER reward has a very nice correlation with the true reward, with the middle-range performance even like a linear correlation. The best model could be selected from 500 and 1000 gradient steps. However, as we train the model for longer, the VIPER reward for the expert trajectories is boosted even higher, and as a side effect, it also relatively boosts up the VIPER reward for low-performance trajectories. This is because, when overfitting the expert trajectories, the model increases the marginal likelihood of all the observations in the expert trajectories to a higher value, which also includes a few frames of the robot lying on the ground at the very beginning of each trajectory after reset. For these low-performance trajectories, the robot remains mainly stuck around the initial position and struggles on the ground. This artifact of the overfitted VIPER reward creates a sharp local maximum in the low-performance region that the agent can hardly get away from.
Appendix H Full results of the benchmark
In this section, we show the full results of all the variants on the 15 tasks on the two benchmark suites. The aggregated results of all the variants are shown in Figure 9 and results on each individual tasks are shown in Figure 10 and Figure 11 respectively for DMC and MetaWorld. We also summarised the final performance after the interaction budget at Table 2 and Table 3 respectively for DMC and MetaWorld.
Figure 9: Comparing different variants of AIME-NoB with other algorithms. The figures show aggregated IQM scores on 9 DMC tasks and 6 MetaWorld tasks. All the algorithms are evaluated with 5 seeds with the shaded region representing 95% CI.